OpenStudy (tt0860):

The length of a rectangle is 3 more than twice its width. The perimeter is 48 feet. What is the length and width of this rectangle? L = 8, W = 5 L = 6, W = 8 L = 17, W = 7 L = 16, W = 8

OpenStudy (_mami_):

still need help??

OpenStudy (_mami_):

@tt0860

OpenStudy (tt0860):

yea

OpenStudy (tt0860):

@_Mami_

OpenStudy (_mami_):

okay so! First.. L=3+2w. p=48 So for length and width of a rectangle: p=2L+2W then.. 2L+3W=48 Make sense so far??

OpenStudy (tt0860):

yes

OpenStudy (_mami_):

2*(3+2W)+2W=48 6+4W+2W=48 6+6W=48 6W=42 W=7